An object moving at 13 m/s slows uniformly with an acceleration of 2.0 m/s2

for a time of 6.0s. Determine a) its final velocity and b) the distance moved in the 6.0 s.


1
Expert's answer
2022-02-28T10:05:38-0500

Given:

"v_0=13\:\rm m/s"

"a=2.0\:\rm m/s^2"

"t=6.0\:\rm s"

a) the final velocity 

"v=v_0+at\\
=13+2.0*6.0=25\:\rm m/s"

b) the distance

"d=v_0t+at^2/2\\
=13*6.0+2.0*6.0^2/2=114\:\rm m"